[Hallc_running] [New Logentry] RC: Daily update

mdiefent at jlab.org mdiefent at jlab.org
Wed Apr 18 09:55:02 EDT 2018

Logentry Text:
<p><strong>Update on day and swing shifts of April 17 and owl shift of April 18</strong></p><p><strong>CEBAF status</strong>:</p><ul><li>RF recovery during the day shift (https://logbooks.jlab.org/entry/3562385). 0.8 hours of beam time. </li><li>RF recovery continues and problems with coming up after RF recovery during the swing shift (https://logbooks.jlab.org/entry/3562592). 2.9 hours of beam time. </li><li>Very stable beam during the owl shift (https://logbooks.jlab.org/entry/3562729). 7.4 hours of beam time. </li><li>Beam studies are tentatively scheduled for Monday, April 23. </li><li>RF recovery is tentatively scheduled for Tuesday, April 24.</li></ul><p><strong>Hall C status</strong>:</p><ul><li>Hall C is configured for the SIDIS experiment (E12-09-017). The HMS is setup for detecting electrons (inclusive measurement) and the SHMS is setup for detecting charged pions and kaons (semi-inclusive measurment). Details on the experimental setup are available on: 
 https://hallcweb.jlab.org/wiki/index.php/E12-09-017_Run_Plan#Experimental_setup </li><li>For various theta angles in the HMS, the beam current is reduced significantly to keep the pTRIG1 rate below 500kHz. </li><li>Controlled access during RF recovery: </li><ul><li>Andrew Kenyon preemptively replaced gearmotors on HMS dipole for helium top fill, bottom fill, and cooldown (https://logbooks.jlab.org/entry/3562366).</li><li>Bill, Brad, and Ryan inspected the gas pressure of the SHMS HGC (https://logbooks.jlab.org/entry/3562286). </li><li>Jack Segal worked on the HMS Dipole power supply (https://logbooks.jlab.org/entry/3562449). He swappd out the AUX PSU for the control board in the power supply and inspected the control board and its wiring. </li><li>Simona inspected the cables for channels 1- to 7+ of the SHMS aerogel and verified that the length of all cables is identical (https://logbooks.jlab.org/entry/3562309). </li><li>Vardan and Vladimir inspected t
 he SHMS aerogel for possible light leaks (ht!
 tps://logbooks.jlab.org/entry/3562401).</li><li>Problems with the sliding door in the SHMS. The roller was removed, the door re-positioned, and roller re-installed (https://logbooks.jlab.org/entry/3562451). </li></ul><li>Studies during the swing shift: </li><ul><li>Harp scans during the swing shift. For 3H07A and 3H07B, the dispersion in y is 1.6 times larger than yesterday’s results while the dispersion in x is unchanged (https://logbooks.jlab.org/entry/3562417). </li><li>Carbon hole centering (https://logbooks.jlab.org/entry/3562473). </li><li>Continued yesterday’s study of the BCM linearity with beam currents of 10muA and 20muA (https://logbooks.jlab.org/entry/3562417). </li></ul><li>Problems with HMS dipole during swing and owl shifts: </li><ul><li>10:36 p.m.: switched to local configuration </li><li>11:11 p.m.: ramped successfully to nominal field </li><li>00:17 p.m.: switched to local configuration </li><li>00:54 a.m.: switched 
 to local configuration while almost ramping up to nominal field </li><li>01:07 a.m., 1:34 a.m., 1:51 a.m.,1:56 a.m.,1:58 a.m.,2:23 a.m.,2:43 a.m.: switched to local configuration while ramping up to nominal field </li><li>During a controlled access, Steve Lassister systematically isolated the problem (based on Jack Segal's and his earlier work). AUX PSU was outputting 4.8V instead of 5.0V. Connectors of cable between AUX PSU and the main board had to be reterminated (https://logbooks.jlab.org/entry/3562615). </li></ul></ul><p><strong>Data taking</strong>:</p><ul><li>The run plan has been updated on April 16 to optimize the physics reach. The detailed plan is available on: https://hallcweb.jlab.org/wiki/index.php/E12-09-017_Run_Plan F</li><li>or kinematic settings 1 and 2, we have completed 10 out of 14 measurements with positive SHMS polarimetry and 12 out of 14 measurements with negative SHMS polarimetry. </li><li>The data taking of the SIDIS experiment with var
 ious configuration changes between the runs !
 requires !
 three persons on shift. We have many open slots in the shift schedule (https://misportal.jlab.org/mis/physics/shiftSchedule/?experimentRunId=HALLC-2017-2018), in particular for the coming weekend and the coming week. We need more volunteers for shifts. </li></ul><p><strong>Hall C Daily RC meeting</strong>: We will meet today at 2:30 p.m. in the conference room located on the second floor of the counting house. Remote participation will be possible via Blue Jeans (Meeting ID: 543903672). You can join via:</p><ul><li>browser: https://bluejeans.com/543903672</li><li>phone: (888) 240-2560 (US toll free), list of international numbers: https://bluejeans.com/numbers</li></ul>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3562809
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180418/3570ec8d/attachment-0001.html>

More information about the Hallc_running mailing list